# On the irreducible character degrees of symmetric groups and their multiplicities

### v1

@@ 3.2 General linear groups
Moreover, there is an addable node to $\lambda$ in the $(4,4)$ position. This gives the partition $\mu=(7^{2},4^{2},3^{2},1)$, with conjugate $\mu^{\prime}=(7,6^{2},4,2^{3})$. As the box lies in the diagonal, $a(\mu)=a(\mu^{\prime})=60$. We can form the partition $\mu^{(m)}$ in analogy with $\lambda^{(m)}$ by adding a box in the $(4,4)$ position to $\lambda^{(m)}$ (or by extending the first row and column of $\mu$) to obtain a partition with the desired property for all odd integers at least $29$.
@@
On the other hand, using a database of all clusters of partitions of size at most 60, we find exactly three pairs of non-conjugate partitions whose unipotent generic degree is the same. The general methods and constructions from [3] give no control over the $a$-function, so this area is wide open. It seems ambitious to offer a conjecture on no evidence at all, so instead we should phrase it as a question.
Question 3.7.
@@
It seems reasonable that the maximal multiplicity will, at some point, exceed $2$, although it does not for $n\leq 84$. (We can check this because all clusters of size at most $84$ are known, and in the symmetric package database.) It is not so clear at this stage whether the generic degree multiplicities should tend to infinity, or even be unbounded.
@@ 4 Largest degrees for symmetric groups
Figure 4.1 gives the ratio $b_{3}(n)/b_{1}(n)$ for $n$ up to $135$, and from this it is easy to see that $b_{3}(n)/b_{1}(n)$ and $b_{2}(n)/b_{1}(n)$ are bounded below by $0.8$ for all $52\leq n\leq 135$, and highly likely that this is true for all larger $n$. If $b_{3}(n)$ and $b_{2}(n)$ are each achieved by a unique (necessarily self-conjugate) character then the sum of the squares of character degrees that are $b_{2}(n)$ or $b_{3}(n)$ is less than $2b_{1}(n)^{2}$, and otherwise it exceeds $2b_{1}(n)^{2}$. (See Conjecture 4.1 below.)
@@
Can both the second- and third-largest degrees be achieved by unique characters? Yes. The cases $n=19,59$ are the only ones for which this occurs up to $n=135$, but there is no reason to believe that it will not occur again. Thus we have the following conjecture, where we have to exclude some small $n$.
@@
There have been a number of papers in the last few decades about large-degree characters, and a continuous version of the hook formula. We specifically mention [22, Theorem 2]; this theorem states, loosely speaking, that the squares of the degrees of the ‘large-degree’ characters of $S_{n}$ sum to $n!-o(n!)$, i.e., most of the order of the group. On the other hand, [22, Theorem 1] bounds $b_{1}(n)$ well away from $\sqrt{n!}$, meaning that there must be unboundedly many large-degree characters of $S_{n}$.
The main barrier to these two theorems implying the following conjecture is that we do not know whether there is only boundedly characters for each of these large degrees $b_{i}(n)$.
@@
Despite several papers being written on the subject (most recently, for example, [7]), we do not have an algorithm to give the largest degree, even conjecturally. Ideas like that the largest character degree should be parametrized by a partition $\lambda$ that is ‘very close’ to a self-conjugate partition (as considered in [23, Section 3] and subsequent papers by this group of authors) appear perhaps to not hold in general, based on our evidence here. For example, if $n=96,109,123$ then one must remove four boxes to arrive at a self-conjugate partition. (These are the only such $n$ up to $135$.) This is despite the fact that $b_{1}(n)$ for $n=94,107,121$ being realized by a self-conjugate partition; in general, one cannot find $b_{1}(n)$ by adding a box to the partition realizing $b_{1}(n-1)$ and choosing the partition with largest degree, i.e., inducing the largest-degree character from $S_{n-1}$. The first $n$ for which one needs to add three boxes to a self-conjugate partition to obtain a partition realizing $b_{1}(n)$ is $n=37$, and then one needs three boxes for $n=45,54,83,95,97,102,108,110,122,124$ up to $135$.
Indeed, one actually needs to subtract five boxes from the partition witnessing $b_{1}(132)$ to obtain a self-conjugate partition. This leads the author to suspect that the answer to the following question is ‘no’.
@@
The question, at first glance, should intuitively have a positive answer. The idea is supposed to be that as $n$ increases one should be able to get arbitrarily close to the LSVK curve, which is self-conjugate, so you should not need to remove more than a couple of boxes to hit something self-conjugate. Unfortunately it does not look from the data as though things are tending towards being closer and closer to a self-conjugate partition, in fact the opposite, given the five boxes needed for $n=132$.
This suggests that any form of ‘shaking algorithm’ [^4] Roughly speaking, a shaking algorithm involves finding a large-degree character of $S_{n}$, restricting it to $S_{n-k}$ for some fixed $k$, then inducing the factors of this back up to $S_{n}$ and looking for one with a larger degree. Combinatorially – using the branching rule – it involves finding all ways to remove $k$ boxes, add $k$ boxes back again, then determine the hook lengths of the resulting partition, hoping for something larger. alluded to above (see, for example, [7]) is doomed to fail to determine $b_{1}(n)$ efficiently in the long run.

### v3

@@ 3.2 General linear groups
Moreover, there is an addable box to $\lambda$ in the $(4,4)$ position. This gives the partition $\mu=(7^{2},4^{2},3^{2},1)$, with conjugate $\mu^{\prime}=(7,6^{2},4,2^{3})$. As the box lies in the diagonal, $a(\mu)=a(\mu^{\prime})=60$. We can form the partition $\mu^{(m)}$ in analogy with $\lambda^{(m)}$ by adding a box in the $(4,4)$ position to $\lambda^{(m)}$ (or by extending the first row and column of $\mu$) to obtain a partition with the desired property for all odd integers at least $29$.
@@
On the other hand, using the database of all clusters of partitions of order at least 4 and size at most 90 in the symmetric package, we find that only 74 of the 13814 such clusters of partitions have unipotent generic degrees that are not all distinct, so making the multiset of hook lengths and the $a$-function coincide is much more restrictive a condition. The general methods and constructions from [3] give no control over the $a$-function, so this area is wide open. It seems ambitious to offer a conjecture on no evidence at all, so instead we should phrase it as a question.
Question 3.8.
@@
It seems reasonable that the maximal multiplicity will, at some point, exceed $2$, although it does not for $n\leq 90$. (We can check this because all clusters of size at most $90$ are known, and in the symmetric package database.) It is not remotely clear at this stage whether the generic degree multiplicities should tend to infinity, or even be unbounded.
@@ 4 Largest degrees for symmetric groups
[…]
@@
Can both the second- and third-largest degrees be achieved by unique characters? Yes. The cases $n=19,59$ are the only ones for which this occurs up to $n=150$, but there is no reason to believe that it will not occur again. Thus we have the following conjecture, where we have to exclude some small $n$.
@@
There have been a number of papers in the last few decades about large-degree characters, and a continuous version of the hook formula. We specifically mention [25, Theorem 2]; this theorem states, loosely speaking, that the squares of the degrees of the ‘large-degree’ characters of $S_{n}$ sum to $n!-o(n!)$, i.e., most of the order of the group. On the other hand, [25, Theorem 1] bounds $b_{1}(n)$ well away from $\sqrt{n!}$, meaning that there must be unboundedly many large-degree characters of $S_{n}$.
The main barrier to these two theorems implying the following conjecture is that we do not know whether there are only boundedly characters for each of these large degrees $b_{i}(n)$.
@@
Despite several papers being written on the subject (most recently, for example, [7]), we do not have an algorithm to give the largest degree, even conjecturally. Ideas like that the largest character degree should be parametrized by a partition $\lambda$ that is ‘very close’ to a self-conjugate partition (as considered in [26, Section 3] and subsequent papers by this group of authors) appear perhaps to not hold in general, based on our evidence here. For example, if $n=96,109,123,138,148$ then one must remove four boxes to arrive at a self-conjugate partition. (These are the only such $n$ up to $150$.) This is despite the fact that $b_{1}(n)$ for $n=94,107,121,136$ being realized by a self-conjugate partition; in general, one cannot find $b_{1}(n)$ by adding a box to the partition realizing $b_{1}(n-1)$ and choosing the partition with largest degree, i.e., inducing the largest-degree character from $S_{n-1}$. (This is not a new observation.) The first $n$ for which one needs to add three boxes to a self-conjugate partition to obtain a partition realizing $b_{1}(n)$ is $n=37$, and then one needs three boxes for $n=45,54,83,95,97,102,108,110,122,124,137,139$ up to $150$.
Indeed, one actually needs to subtract five boxes from the partition witnessing $b_{1}(132)$ to obtain a self-conjugate partition, and then again for $n=147,149$. This leads the author to suspect that the answer to the following question is ‘no’.
@@
The reason that this question is relevant is that the methods and algorithms to generate large-degree characters, for example given in [7], are dependent on starting with something close to $b_{1}(n-1)$ and obtaining something close to $b_{1}(n)$ from it by adding a box. The behaviour of the function that sends $n$ to the number of boxes needed to be added to a self-conjugate partition to obtain partition with character degree $b_{1}(n)$ is that it is not smooth, and frequently jumps around, by ever-increasing amounts. This suggests that any form of ‘shaking algorithm’ [^8] Roughly speaking, a shaking algorithm involves finding a large-degree character of $S_{n}$, restricting it to $S_{n-k}$ for some fixed $k$, then inducing the factors of this back up to $S_{n}$ and looking for one with a larger degree. Combinatorially – using the branching rule – it involves finding all ways to remove $k$ boxes, add $k$ boxes back again, then determine the hook lengths of the resulting partition, hoping for something larger. alluded to above (see, for example, [7]) is doomed to fail to determine $b_{1}(n)$ efficiently in the long run.
On the other hand, $b_{1}(n)$ is achieved by a self-conjugate partition quite often – for larger $n$, between $120$ and $150$ it is achieved for $n=121,125,127,129,134,136,142,144$. It is highly likely that $b_{1}(n)$ is realized by a self-conjugate partition infinitely often, but the author is not aware of any progress on that question. Finding a pattern to the $n$ here seems non-obvious. [^9] I would like to thank one of the referees for suggesting these questions in this section.